OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: 2x – y > 7 5x – y > 2

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=2x%20%E2%80%93%20y%20%3E%207%2C%205x%20%E2%80%93%20y%20%3E%202&t=mfftb01 choose any point on that shaded region, that is your answer.

OpenStudy (turingtest):

There are infinite solutions, but (5,0) would work because 2(5)-0=10>7 5(5)-0=25>2 @experimentX stop using Wolfram for everything, most people here need human explanations!!!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you can let x=6 and y=3 2(6)-3=9>7

OpenStudy (anonymous):

x=5 and any number higher and y=0 will work also

OpenStudy (anonymous):

excuse me I mean 4